OpenStudy (anonymous):

What percent of 67 is 33? Round to the nearest tenth of a percent

OpenStudy (austinl):

\(\displaystyle \frac{x}{100}=\frac{33}{67}\) Solve for x!

OpenStudy (anonymous):

100*33/67 = 49.254 = 49.3%?

OpenStudy (austinl):

Yep, good job!

OpenStudy (anonymous):

thanks :)